Gold Creek Fire Update 07/25/2013
Incident: Gold Creek Wildfire
A smaller Type III organization took command of the fire at 6:00 a.m. today.
A morning update will continue to be published as long as fire conditions require.
Start date: Identified July 15 Contact: Roger Peterson
Phone: 406-396-9216 Email: firstname.lastname@example.org
Location: 16 miles northeast of Missoula, MT.
Current Size: Approximately 170 acres Containment: 98%
Resources Assigned: 1 Type I handcrew, 2 Type II handcrews, 1 Wildland Fire Module, 2 Helicopters. Total personnel assigned are 208.
Transition: Crews will patrol through the end of shift today. Transition of operations to a Type 3 team occurred at 6 a.m. this morning.
Closures: Gold Creek Area Closure – Gold Creek Trailhead and West Fork Gold Creek Trailhead will be closed. The area closure extends to all areas west of Gold Creek Road, north of Trail #52 (Primm Lake Trail) and east of the Main Rattlesnake Recreation Corridor. Trail #504 (Porcupine Ridge Trail) is closed. Trail #52 and trails south of the Rattlesnake vicinity are open. The closure also borders Trail #515 accessing the headwaters of Rattlesnake Creek, this trail is open. Closure map located on InciWeb.
Forest Road #126 will remain closed at Cow Creek Gate (approx. 7 miles in), the West Fork of Gold Creek Trailhead (Rd #4323) at the junction of Rd #4323 and Rd #2103 is currently closed. This is approx. 6 miles to West Fork of Gold Creek Trailhead.
Current Status: Yesterday was a good day with continued mop-up and gridding. Crews in Division A along the western flank reported no smoke. Division B worked the northwest flank and gridded for spots and continued mop-up operations extinguishing hot spots from a handful of smokes that will be an area of focus today. Crews in Division E along the southeast flank conducted extensive gridding along the fire’s edge searching for spots and debris, successfully mopping up a large spot fire. One smoke yesterday morning in Division O along the southwestern corner of the fire was extinguished. Crews continued to search the green and grid the interior of the fire.
Planned Actions: Crews will continue to mop up all flanks of the fire by searching for hot spots and digging up the burning material using water where necessary and extinguishing embers. Patrols will continue in all divisions. A Type 3 camp has been established at Drop Point 1. Transition of operations to a Type 3 team occurred at 6 a.m. today.
Weather: Thursday will be another hot and dry day temperatures are forecast in the mid-80’s with relative humidity between 15-17% and steady light winds with gusts up to 10 m.p.h. Be alert to possible scattered thunderstorms cells in the afternoon through Friday. Winds will increase on Saturday.
Cooperating Agencies: Confederated Salish and Kootenai Tribes, Montana DNRC, Lolo National Forest.
The fire will be listed on Inciweb at http://inciweb.nwcg.gov/.